OpenStudy (anonymous):

Find f. f ''(x) = 20x^3 + 12x^2 + 4 f(0) = 3 f(1) = 3 My Answer was x^5+x^4-2x+3 my answer was wrong, can someone tell me where I went wrong?

OpenStudy (anonymous):

one sec

myininaya (myininaya):

f'(x)=20x^4/4+12x^3/3+4x+C=5x^4+4x^3+4x+C f(x)=5x^5/5+4x^4/4+4x^2/2+Cx+D=x^5+x^4+2x^2+Cx+D

OpenStudy (anonymous):

what math are you in Calc I or II

OpenStudy (anonymous):

1

OpenStudy (anonymous):

okay I used integration, I'm assuming that you know how?

OpenStudy (anonymous):

Yes

myininaya (myininaya):

f(0)=D=3 f(1)=1+1+2+C+3=3 so C=-4

OpenStudy (anonymous):

okay so take the integral of f double prime (c)

OpenStudy (anonymous):

i mean (x)

OpenStudy (anonymous):

so you get 20x^4/4 + 12x^3/3 + 4x

OpenStudy (anonymous):

then simplify and do it once more

OpenStudy (anonymous):

go it?

OpenStudy (anonymous):

Medal pleaser ^.^

OpenStudy (anonymous):

f(x)= x^5+x^4+C(x)+D

OpenStudy (anonymous):

Plug f(0) in for x and set it equal to 3 makes D= 3 Plug f(1) in for x and set it equal to 3 makes C= -2 plugging those in was wrong

myininaya (myininaya):

you are missing 2x^2 in f(x)

OpenStudy (anonymous):

x^5 + x^4 + 2x^2 + c

OpenStudy (anonymous):

myininaya solved it already

OpenStudy (anonymous):

yeah

myininaya (myininaya):

f(x)=x^5+x^4+2x^2+Cx+D

OpenStudy (anonymous):

Thanks all